OpenStudy (anonymous):

write 50/8 as a reduced mixed number

OpenStudy (mathstudent55):

50/8 = 25/4 = 6 remainder 1 = 6 1/4

OpenStudy (anonymous):

ultiplying. 1/2 X 2/3 multiply the numerators together, then multiply the denominators. 1X2=2 2X3=6 the answer is 2/6 reduced= 1/3 DIVIDING FRACTIONS * Invert (i.e. turn over) the denominator fraction and multiply the fractions * Multiply the numerators of the fractions * Multiply the denominators of the fractions * Place the product of the numerators over the product of the denominators * Simplify the Fraction Example: Divide 2/9 and 3/12 * Invert the denominator fraction and multiply (2/9 ÷ 3/12 = 2/9 * 12/3) * Multiply the numerators (2*12=24) * Multiply the denominators (9*3=27) * Place the product of the numerators over the product of the denominators (24/27) * Simplify the Fraction (24/27 = 8/9) * * The Easy Way. After inverting, it is often simplest to "cancel" before doing the multiplication. Canceling is dividing one factor of the numerator and one factor of the denominator by the same number. * For example: 2/9 ÷ 3/12 = 2/9*12/3 = (2*12)/(9*3) = (2*4)/(3*3) = 8/9 ~~~~~~~~~~~~~~~~~~~~~~~ adding and subtracting mixed numbers. 1 2/3 + 3 1/3 multiply the denominator and the whole number together than add the . 3X1=3+2=5 5/3 3X3=9+1=10 10/3 so its 5/3 + 10/3 15/3 the answer is 5 <--whole number subtracting is the same, you just subtract. multiplying mixed numbers. make the mixed number an improper fraction like i did above. then multiply across just like you do for regular multiplying. 1 2/3 X 3 2/3 3X1=3+2=5= 5/3 3X3=9+2=11=11/3 11X5=55 3X3=9 answer= 55/9 covert that into a mixed number. 55/9=6 9X6=54 55-54=1 answer= 6 1/9 dividing mixed numbers Dividing two mixed numbers: Convert each mixed number to an improper fraction. Invert the improper fraction that is the divisor. Multiply the two numerators together. Multiply the two denominators together. Convert the result back to a mixed number if it is an improper fraction. Simplify the mixed number. Example: 6 2/8 ÷ 3 5/9 = Convert each mixed number to an improper fraction. 50/8 ÷ 32/9 Invert the improper fraction that is the divisor and multiply. 50/8 * 9/32 Multiply the two numerators together. 50 * 9 = 450 Multiply the two denominators together. 8 * 32 = 256 Convert the result back to a mixed number. 450/256 = 1 194/256 Simplify the mixed number. 1 97/128 i hope this helps :)